OpenStudy (anonymous):

Use the compound interest formula, Alexia has $4864.77 in her money market account currently. If the annual interest rate is 3.9%, and the interest is compounded twice per year, how much money will be in her account after 2 years? A. $4,149.57 B. $5,244.22 C. $5,255.47 D. $15,733.31

OpenStudy (anonymous):

@texaschic101 can you help me again??

OpenStudy (anonymous):

@noneyabusiness can you help me??

OpenStudy (anonymous):

@uri can you help me??

OpenStudy (anonymous):

@Xan can you help me??

OpenStudy (anonymous):

@Pianolover03

OpenStudy (anonymous):

Your equation is \[A = 4864.77(1 + \frac{ 0.039 }{ 2 })\ ^{\left( 2 \times 2 \right)}\] Can you solve it?

OpenStudy (anonymous):

\[A = 4864.77(1 + \frac{ 0.039 }{ 2 })^{\left( 2 \times 2 \right)}\] \[A = 4864.77(1 + 0.0195)^{\left( 4\right)}\] \[A = 4864.77(1.0195)^{\left( 4\right)}\] \[A = 4864.77(1.0803113)\] \[A = 5255.466\] \[A = 5255.47\] So your answer is C.